# Revisiting Monocular 3D Object Detection from Scene-Level Depth Retargeting to Instance-Level Spatial Refinement

### v1

@@ 4.4 Ablations
To illustrate the effectiveness of SDR and ISR modules proposed in this paper, we carry out ablations on both KITTI test and val sets as shown in Table 3. We compare the performance on Car category under IoU criterion 0.7.
SDR: Exp. (a) is the baseline which directly uses depth one-hot encoding to generate 3D features for 3D object detection. In Exp. (b), when retargeting depth representation from depth one-hot encoding to depth thickness field, the detection $AP_{3D}$ increases from 26.95/16.15/13.45 to 31.12/20.04/17.33 on the test set. ISR: As shown in Exp. (c), when enforcing the refinement of voxel spatial representation via learning 3D occupation with the guidance of instances, the detection $AP_{3D}$ is boosted to 31.97/19.79/17.38 on the test set. Finally, we can see RD3D obtains +5.25/+5.04/+5.35 in $AP_{3D}$ compared to the baseline from Exp. (d).
Universality Analysis. In Table 4, we report the performances of RD3D when equipped with different depth estimation models. DAMv2 and Metric3Dv2 are two large models for depth estimation, showing significant improvements in detection performance when integrated with our method. We have fine-tuned DAMv2 which was initially designed as a relative depth model on KITTI dataset for absolute depth estimation. Its results demonstrate that our method is well-suited to the current paradigm in depth estimation models: initial self-supervised training on relative depth, followed by fine-tuning for absolute depth. As an early lightweight model for depth estimation, Monodepth2 still benefits from our method for monocular 3D object detection. These experiments demonstrate both the effectiveness and the universality of RD3D.
vs. Traditional Depth Representations. In Table 5, we report the performance when replacing the depth thickness field with traditional depth representations. We can observe that the depth thickness field achieves better performance compared to depth one-hot encoding and depth distribution.

### v2

@@ 4.4 Ablations
To illustrate the effectiveness of SDR and ISR modules proposed in this paper, we carry out ablations on both KITTI test and val sets as shown in Table 3. SDR: Exp. (a) is the baseline which directly uses depth one-hot encoding to generate 3D features for 3D object detection. In Exp. (b), when retargeting depth representation from depth one-hot encoding to depth thickness field, the detection $AP_{3D}$ increases from 26.95/16.15/13.45 to 31.12/20.04/17.33 on the test set. ISR: As shown in Exp. (c), when enforcing the refinement of voxel spatial representation via learning 3D occupation with the guidance of instances, the detection $AP_{3D}$ is boosted to 31.97/19.79/17.38 on the test set. Finally, we can see that MonoDTF obtains +5.25/+5.04/+5.35 in $AP_{3D}$ compared to the baseline of Exp. (d).
Universality Analysis. In Table 4, we report the performance of MonoDTF when equipped with different depth estimation models. DAMv2 and Metric3Dv2 are large models used for depth estimation, distinguished by their methodology. DAMv2 initially produces relative depth, which requires fine-tuning of the metric depth data for 3D detection. In contrast, Metric3Dv2 provides metric depth directly, obviating the requirement for fine-tuning. The results indicate significant improvements in detection performance when integrated with our method. As an early lightweight model for depth estimation, Monodepth2 still benefits from our method for monocular 3D object detection. These experiments demonstrate both the effectiveness and the universality of MonoDTF.
Ablations on Depth Representations. In Table 5, we report the performances when replacing the depth thickness field with traditional depth representations (e.g., depth one-hot encoding and depth distribution). We can observe that the depth thickness field performs better than the traditional depth representations.